# Quantifying the Reliability of Predictions in Detection Transformers: Object-Level Calibration and Image-Level Uncertainty

### v2

@@ III-C Expected Calibration Error (ECE) Metrics
Following the standard protocol [20], two IoU thresholds ($\tau=0.5$ and $\tau=0.75$) are used, and the D-ECE scores obtained under these thresholds are averaged to yield the final metric. Because this process is carried out on a validation set, it provides an overall measure of the model’s calibration performance on that set. A lower D-ECE indicates that the confidence scores are better aligned with actual precision, while a higher D-ECE reveals a potential mismatch between confidence and true detection performance.
LaECE. To account for both localization error and classification precision, localisation-aware expected calibration error (LaECE) [19] aligns model confidence with the product of the predicted bounding box’s precision and its IoU with the ground-truth box. Specifically, LaECE is computed in a class-wise manner to mitigate class imbalance. Let $c$ index the class, and let $\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$. Partition the confidence scores into $J$ bins, and let $\widehat{\mathcal{D}}_{j}^{c}\subseteq\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$ in bin $j$. Define $\bar{p}^{c}_{j}$ to be the average confidence in $\widehat{\mathcal{D}}_{j}^{c}$, $\mathrm{precision}^{c}(j)$ to be the corresponding precision, and $\bar{\mathrm{IoU}}^{c}(j)$ to be the average IoU of $\widehat{\mathcal{D}}_{j}^{c}$. The LaECE metric is then given by:
$\text{LaECE}\triangleq\frac{1}{K}\sum_{c=1}^{K}\sum_{j=1}^{J}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}_{j}^{c}-\text{precision}^{c}(j)\times\bar{\text{IoU}}^{c}(j)\Bigr{|}$ (3)
where $K$ is the total number of classes, and an IoU threshold $\tau$ is employed to determine whether a detection is considered a true positive.
LaECE0 [20] sets the IoU threshold $\tau$ to zero and assigns an average IoU of zero to false positive detections, thereby reducing the original formulation. In particular, if a detection is not a true positive under $\tau=0$, its IoU contribution is taken to be zero. The resulting metric becomes:
$\text{LaECE}_{0}=\frac{1}{K}\sum^{K}_{c=1}\sum^{J}_{j=1}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}^{c}_{j}-\bar{\text{IoU}}^{c}(j)\Bigr{|}.$ (4)
By setting $\tau=0$, false positives have $\bar{\mathrm{IoU}}=0$, and thus the calibration criterion focuses more explicitly on the contribution of bounding box localization to the overall calibration error.
III-D Image-Level Reliability
@@
We define image-level reliability as a measure of how accurately and confidently the predictions match the ground truth objects:
$\mathsf{ImReli}(\bm{x};\theta)\triangleq\mathsf{Perf}\big{(}\hat{\mathcal{D}}_{\theta}(\bm{x}),~{}\mathcal{D}_{\bm{x}}\big{)}$ (5)
@@
where one can use any standard performance metrics, such as average precision and recall, for $\mathsf{Perf}$ depending on the user’s requirements.

### v3

@@ V-C Numerical Analysis
Optimal Positives D-ECE 1.000 0.983 0.791 1.000 0.963 0.967 1.000 0.890 0.989 0.954 $\pm$ 0.066
[…]
$\mathsf{OCE}^{*}$ 0.791 0.859 1.000 0.967 0.873 1.000 0.989 0.854 1.000 0.926 $\pm$ 0.076
V-C1 Validating OCE as a Calibration Metric
Our first step is to validate that OCE is a rightful measure of calibration. We first perform a sanity check to ensure OCE fundamentally measures the same concept of calibration as established metrics. To do this, we isolate the metrics’ behaviors from the confounding effects of different post-processing decisions. We therefore choose the optimal positive set. This set best represents the predictions that each model was specifically trained to produce as a calibrated predictor.
Hypothesis. If OCE is a valid calibration metric, its ranking of models on this oracle optimal positive set should strongly correlate with the rankings produced by established metrics (D-ECE, LA-ECE0). A lack of correlation would suggest OCE is measuring a completely different, unrelated property.
Method. We evaluate the calibration errors on the optimal positive set from our six models—UP-DETR, D-DETR, Cal-DETR, DINO, Faster-RCNN, and YOLOv3—using OCE, D-ECE, and LA-ECE0. We then evaluate the models using each metric and compute the Pearson correlation coefficient (PCC) between metrics.
As shown in the first three rows in Table 1, the correlations are all extremely high. This robust finding provides a critical validation: OCE captures the same underlying concept of calibration quality as established metrics. This confirms OCE is a legitimate calibration metric, providing the necessary foundation for the next section, where we will demonstrate its superiority in the practical scenario where optimal positive set is not accessible.
V-C2 Pitfalls of Existing Metrics as a Reliable Set Selection Criteria
We now analyze the behavior of existing metrics and the proposed OCE in a practical scenario: selecting a confidence threshold (i.e., post-processing configuration) that achieves the lowest calibration error for a given model.
To this end, this section analyzes the effectiveness of existing calibration metrics (D-ECE, LA-ECE0, and our OCE) as well as performance metrics (AP [16, 17, 18] and Localization recall precision (LRP) [21, 36]) for this task. Specifically, for each of the four different DETR models, we sweep the confidence score threshold from $0.0$ to $1.0$. This generates different positive subsets of predictions. We then plot the score of each metric for these subsets to determine the threshold that each metric prefers (i.e., the one that yields its optimal score). Results on Cal-DETR are illustrated in Figure 8. For definitions of metrics and additional results on other DETR models, please refer to the Appendix C. Our empirical findings reveal the following behaviors:
Pitfall 1: Average Precision Prefer No Filtering. As noted in several studies [21, 19, 20, 36], the optimal AP is achieved when the threshold is set close to $0.0$. This is because AP does not harshly penalize low-confidence predictions. However, as discussed so far, using the full prediction set (hundreds of predictions) is practically non-sensical. It carries a high risk of including uncalibrated negative predictions, diminishes interpretability, and leads to unreliable downstream decisions.
Pitfall 2: ECEs Prefer Discarding Almost All Predictions. More importantly, the optimal ECEs (D-ECE and LA-ECE0) are achieved when the threshold is set close to $1.0$. This behavior, which favors retaining only a tiny subset of high-confidence predictions, is a structural pitfall of prediction-level ECEs [20]. The cause is simple: ECEs do not penalize missed predictions (i.e., false negatives). As a result, they can achieve a near-zero (perfect) error by evaluating a set containing only a few highly accurate and confident predictions, while ignoring all the ground-truth objects the model missed at that high threshold.
Therefore, this pitfall makes ECEs fundamentally unsuitable for the goal we defined: finding a reliable (model + post-processing) combination. If a practitioner uses D-ECE to select a threshold, the metric will misleadingly guide them to select a threshold near 1.0, which is nonsensical as it discards almost all predictions to achieve a low error.
A Note on LRP: Good for Localization, Not Calibration. On the other hand, LRP, a localization-focused performance metric, does appear to find a sensible, non-extreme optimal threshold. This is rational: optimal positives are trained to be well-localized predictors, while optimal negatives are relatively not. LRP, by focusing purely on localization quality, thus correctly identifies a threshold that filters out the poorly-localized predictions. However, LRP is not designed to measure calibration and does not consider the confidence quality of the predictions at all. As a result, a model ranking scored by LRP reflects localization performance, which is not necessarily aligned with the model’s confidence calibration qualities.
V-C3 The effectiveness of OCE
The biggest advantage of using OCE over other metrics is that OCE is capable of measuring the model’s calibration quality alongside the post-processing scheme employed. Resultingly, as illustrated in Figures 8 and 14, the confidence threshold vs. OCE curve is U-shaped and the optimal OCE is achieved around a threshold of $0.3$. This is particularly notable because it reflects the practical choices in many calibration studies to date [8, 9, 10, 11, 12, 13], confirming the reliability of using OCE.
Using a fixed separation threshold of $0.3$ might serve as a reasonable approximation for distinguishing between optimal positives and negatives. However, trained or post-calibrated DETR models often have varying optimal confidence thresholds. For example, UP-DETR has an optimal threshold of approximately $0.5$, and low-temperature calibrated models are likely to require even higher thresholds. Consequently, employing a fixed threshold poses potential risks.
(a) Threshold Optimizing AP
(b) Threshold Optimizing OCE
(c) Threshold Optimizing D-ECE
Figure 9: Exemplary visualization demonstrating the impact of confidence threshold ($\rho$) selection on the final prediction subset in Cal-DETR across different post-processing configurations. Optimal positive and negative predictions are highlighted with blue and red boxes, respectively. The threshold selected by AP (e.g., $\rho=0.1$) retains redundant predictions, mixing calibrated positives with uncalibrated negatives. The threshold selected by D-ECE (e.g., $\rho=0.9$) discards nearly all detections, including the optimal positive prediction with moderate confidence, resulting in a missed detection. The proposed OCE identifies a practical threshold (e.g., $\rho=0.27$) that balances precision and recall while maintaining calibrated predictions.
Figure 10: A visualization of the difference in calibration between optimal positive and negative predictions on the COCO dataset. Object-level Calibration Error (OCE) value of each set is shown, where a lower score represents better-calibrated predictions. As shown, OCE varies significantly between optimal positive and negative predictions, indicating that OCE is a useful metric for assessing the reliability of post-processing schemes.
Figure 11: Pearson Correlation Coefficient (PCC) between per-image AP scores and per-image average foreground confidence scores on optimal positive and negative predictions (higher values indicate stronger correlation). As expected, optimal positive predictions show high positive correlation between confidence and performance. More interestingly, optimal negative predictions exhibit an inverse correlation with image-level reliability.
VI Quantifying Image-level Reliability
In this section, we revisit the importance of separation between primary and secondary predictions at test-time, now from the perspective of image-level reliability estimation.
Recall that our theoretical analysis (Proposition 3) shows that when the model is uncertain about its prediction—expressed by matching probability $P_{i}<1$—it suppresses the predictive distribution. Reflecting this observation to the secondary prediction, we find that its foreground confidence score is no longer zero. Consequently, we expect the separation between these predictions to become blurred under uncertainty, and our empirical observations confirm this.
Specifically, for reliable images, we observe that the confidence of optimal positive predictions increases across decoder layers while optimal negative predictions remain low, resulting in a large gap between optimal positive and negative predictions (e.g., Figure 4a). In contrast, for unreliable instances, optimal positive prediction confidence fails to increase across layers, while optimal negative prediction scores are either slightly elevated or unchanged. This produces a small gap between optimal positive and negative predictions (e.g., Figure 4b).
To further empirically verify this, we first obtain the image-wise AP score by using the same COCO evaluator [18], passing the predictions and annotations for each image individually rather than the entire image set. We use this value as a proxy of image-level reliability (See Definition 2 for more formal interpretation.) Then, we measure the Pearson correlation coefficient between the image-wise AP and the average confidence score of the optimal positives and the optimal negatives, respectively.
As shown in Figure 11, the average confidence scores of the optimal positive predictions exhibit a moderately strong positive correlation with image-level reliability (i.e., the model’s actual average precision on that image) across different models. Notably, the average confidence scores for negative predictions are inversely correlated with image-level reliability. This finding reinforces our claim, highlighting the importance of identifying primary and secondary predictions, when conducting image-level uncertainty quantification (UQ) in DETR.
VI-A Image-Level Reliability
@@
We define image-level reliability as a measure of how closely the predictions align with the ground-truth annotations for a given test image $x$:
$\mathsf{ImReli}(x;\theta)\triangleq\mathsf{Perf}\big(\hat{\mathcal{Y}}_{\theta}(x),~\mathcal{Y}\big),$
@@
where one can use any standard performance metric, such as average precision, for $\mathsf{Perf}$ depending on the user’s requirements.